find the quotient -11 over 24 divide 7 over 10

Found 2 solutions by checkley77, rfer:
Answer by checkley77(12844)

-11 over 24 divide 7 over 10
(-11/24)/(7/10) invert the second fraction & multiply.
-11/24*10/7
-110/168=-55/84 ans.

Answer by rfer(16322)

(-11/24)/(7/10)
-11/24*10/7=-110/168=-55/84